OpenStudy (anonymous):

this experssion represents the volume of a sphere : x^3-8x^2+18x if x =4,what is the volume of the sphere show your work.

OpenStudy (anonymous):

Just plug in 4 for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

put in 4^3-8*4^2+18*4 and hit the enter or = sign

OpenStudy (anonymous):

what doy you get

OpenStudy (anonymous):

i got 42 cubic units

OpenStudy (anonymous):

You should get 8 did you type it in exactly like above... do you have the ^ key?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

how?/////////////////////////////////

OpenStudy (anonymous):

4^3 = 64 -8^4^2 = -128 18*4 = 72 64 - 128 + 72 = 8